OpenStudy (anonymous):

x squared - 10x +? complete the square

OpenStudy (across):

When you are faced with the problem \[x^2+cx+?,\]finding the term that "completes the square" follows:\[x^2+cx+\frac{c^2}{4}=\left(x+\frac{c}{2}\right)^2.\]

OpenStudy (anonymous):

so is it 25

OpenStudy (across):

That's correct. Always try it out and see for yourself:\[x^2-10x+25=(x-5)^2.\]